Light enters train air to glass having refractive index 1.50. What is the speed of light in the glass? The speed ref light in vacuum is 3 × 108 m/s.

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

The speed of light in vacuum is = 3 × 108m/s
Refractive index of glass ng = 1.50
ng = c/v
speed of light in glass vg=c/ ng= 3 x 108
1.50
= 2 x 108m/s

Thanks